OpenStudy (anonymous):

what is the product of (x-7)(2x^2-5x+3)

OpenStudy (anonymous):

someone help please

OpenStudy (mathmale):

Let's walk thru the basic processes required: What is (x)(2x^2)?

OpenStudy (mathmale):

\[x*2x^2=?\]

OpenStudy (anonymous):

what do you mean? like graphed?

OpenStudy (mathmale):

No graphing necessary. You're doing algebraic multiplication. Examples:\[x*x=x^2; ~ x^1*x^1=x ^{1+1}=x^2;~ x^5*x^3=x ^{5+3}=x^8.\]

OpenStudy (mathmale):

Have you seen this algebraic multiplication before?

OpenStudy (anonymous):

i don't know how to do this

OpenStudy (jdoe0001):

\(\large { ({\color{red}{ x}}{\color{blue}{ -7}})(2x^2-5x+3) \\ \quad \\ \quad \\ \begin{array}{cllll} 2x^2-5x+3\\ \times {\color{red}{ x}}\\ \hline \end{array}+ \begin{array}{cllll} 2x^2-5x+3\\ \times {\color{blue}{ -7}}\\ \hline \end{array}\implies product }\)

OpenStudy (mathmale):

For now, please copy down the examples I've given you. What I've done in those examples entails applying "rules of exponents." Each quantity I've given you has x to some power in it, for example, \[x^1;~x^2;~x^3;\]What is the next higher power of x?

OpenStudy (anonymous):

can you just give me the answer please

OpenStudy (mathmale):

I'm sorry you're frustrated! But one of the rules we agree to before we use OpenStudy is that we neither ask for answers outright nor provide them. Still, the next power of x is x^4. What's the next one?

OpenStudy (anonymous):

my spiral's do in 15 minutes and i need answers and x^4 power is 4x right?

OpenStudy (mathmale):

\[x^1;~x^2;~x^3;~x^4;~x^?\]

OpenStudy (mathmale):

Waye, I'd love to help you. But I'm not simply giving out answers because your homework is due in a few minutes. Either now, or later on, let's go through the steps necessary to multiply powers of x together. Your choice.

OpenStudy (mathmale):

What course are you in? Haven't you seen examples of this multiplication of powers of x before?

OpenStudy (mathmale):

Waye: I don't know your circumstances; I'm trying to determine what they are. I do ask that you come back on OpenStudy at least hours, or even days, before your problem set is due, so that we have adequate time to work through the material.

OpenStudy (anonymous):

yeah but i get distracted in class and its hard to pay attention

OpenStudy (mathmale):

I can't help a lot with that; obviously it's up to you to pay attention, ask questions and not allow yourself to become distracted. Once again, Waye, I'd be happy to help you IF you have paid attention, read (or listened to) some examples, and tried doing a couple of problems yourself. If I (or anyone else on OpenStudy) were to just give you answers, you'd not learn very much, or remember anything.

OpenStudy (mathmale):

Come back, Waye, but before you do that, PLEASE do some advance preparation.

OpenStudy (anonymous):

thanks failed my spiral

OpenStudy (mathmale):

Come back when you're up to it and we'll work on these problems. Try to set aside adequate time to study new topics, too.